# Nonmonotone subgradient methods based on a local descent lemma

### v1

@@ Abstract
The aim of this paper is to extend the context of nonmonotone descent methods to the class of nonsmooth and nonconvex functions called upper-$\mathcal{C}^{2}$, which satisfy a nonsmooth and local version of the descent lemma. Under this assumption, we propose a general subgradient method that performs a nonmonotone linesearch, and we prove subsequential convergence to a stationary point of the optimization problem. Our approach allows us to cover the setting of various subgradient algorithms, including Newton and quasi-Newton methods. In addition, we propose a specification of the general scheme, named Self-adaptive Nonmonotone Subgradient Method (SNSM), which automatically updates the parameters of the linesearch. Particular attention is paid to the minimum sum-of-squares clustering problem, for which we provide a concrete implementation of SNSM. We conclude with some numerical experiments where we exhibit the advantages of SNSM in comparison with some known algorithms.
@@ 1 Introduction
where the objective function $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ is assumed to be upper-$\mathcal{C}^{2}$. This is a class of functions that satisfy a nonsmooth local version of the descent lemma, a property that makes them suitable for extending the framework of linesearch methods (for details, see Section 3). This is the main idea that we aim to convey in this work, which is not only of theoretical interest but also practical, since upper-$\mathcal{C}^{2}$ functions naturally arise in different applications, as data mining.
Let us recall that descent methods are a class of iterative optimization algorithms designed to find (local) minima of differentiable functions. Their core technique consists in updating the current iterate by moving in a direction that reduces the value of the objective function. Hence, given a starting point $x_{0}\in\mathbb{R}^{n}$, a general iteration step is defined by
@@
Although nonmonotone line searches bring opportunities to improve the performance of optimization algorithms, they also require the tuning of additional parameters, such as the linesearch stepsize (which is initially chosen at each iteration) and the memory parameter (which controls how many past function values are taken into account). A wrong adjustment of these parameters may slow down the resulting algorithm, and one may lose all possible benefits of nonmonotone line searches. With this possible weakness in mind, we propose a specification of the main scheme, that we name the Self-adaptive Nonmonotone Subgradient Method (SNSM), which incorporates an automatic procedure for selecting the stepsize and memory parameters of the linesearch.
@@ 2 Preliminaries and notation
A function $f:\mathbb{R}^{n}\to\mathbb{R}$ is said to be of class $\mathcal{C}^{k}$ if it is differentiable and all its partial derivatives up to order $k$ are continuous. In particular, we say that a function $f$ is smooth if it is of class $\mathcal{C}^{1}$, i.e., it is differentiable with continuous gradient. If, in addition, its gradient is $L$-Lipschitz continuous we say that $f$ is $L$-smooth. The class of differentiable functions with locally Lipschitz continuous gradient is denoted by $\mathcal{C}^{1,+}$. We represent the gradient and Hessian of $f$ as $\nabla f$ and $\nabla^{2}f$, respectively.
@@ 3 Upper-$\mathcal{C}^{2}$ functions and the descent lemma
Most known gradient-based algorithms are designed to ensure descent of the objective function per iteration. Many of the algorithms proposed in the literature that rely exclusively on first- and second-order information are developed for the setting of differentiable functions with at least locally Lipschitz continuous gradients. This class of functions is particularly important in both theoretical and practical optimization. Its significance lies in the fact that such functions satisfy the so-called descent lemma (see, e.g., [23, Lemma A.11]). More precisely, for a differentiable function $f:\mathbb{R}^{n}\to\mathbb{R}$ with an $L_{f}$-Lipschitz continuous gradient, it holds that
$f(y)\leq f(x)+\langle\nabla f(x),y-x\rangle+\frac{L_{f}}{2}\|y-x\|^{2},\quad\text{for all }x,y\in\mathbb{R}^{n}.$ (1)
@@
The work [6] investigated objective functions that can be expressed as the difference of two functions, namely $\varphi:=g-h$, where $g$ is differentiable with locally Lipschitz continuous gradient, and $h$ is a locally Lipschitz prox-regular function (see, e.g., [33, Definition 13.27]). The authors developed a descent method for the objective function $\varphi$, where descent directions are generated using coderivative information. Among the results presented, they proved that this class of functions satisfies a local version of the descent lemma (see [6, Lemma 3.6]).
@@
Let us recall a related definition extracted from the book of Rockafellar and Wets (see [33, Definition 10.29]) that will serve us to encompass both frameworks [6] and [8]. It refers to a class of functions that is usually known as lower-$\mathcal{C}^{2}$ functions (correspondingly, a function $f$ is called upper-$\mathcal{C}^{2}$ if f-f is lower-$\mathcal{C}^{2}$). The properties of these functions have been well-studied in the context of variational analysis (see, e.g., [12, 29] and the references therein). However, to the best of our knowledge, this class has not received much attention in the numerical optimization literature. Beyond the already mentioned exceptions, we are only aware of [38], where a sequential quadratic programming trust-region method is proposed for the problem of minimizing a Lipschitz continuous and upper-$\mathcal{C}^{2}$ function in a bounded region including $L$-smooth equality constraints.
@@
It is easy to see that a function $\varphi$ is upper-$\mathcal{C}^{2}$ on an open set $U$ if and only if it is upper-$\mathcal{C}^{2}$ around any point $x\in U$. Moreover, by [33, Theorem 10.31], these functions are locally Lipschitz. The next result summarizes some useful characterizations of upper-$\mathcal{C}^{2}$ functions.
@@
where $g(x):=\kappa\|x\|^{2}$ and $h(x):=\max_{c\in C}\{\langle a(c),x\rangle+b(c)\}$. This shows the implication (ii) $\Rightarrow$ (iv). Now, let us suppose that there exists some neighborhood $V$ of $\bar{x}$ and a representation $\varphi=g-h$, where $g$ is differentiable with Lipschitz gradient on $V$ and $h$ is Lipschitz continuous and prox-regular on $V$. It follows from [33, Proposition 13.34] that $g$ is upper-$\mathcal{C}^{2}$ on $V$, and by [33, Proposition 13.33], the function h-h is upper-$\mathcal{C}^{2}$ on $V$. This concludes the proof, since the class of upper-$\mathcal{C}^{2}$ functions is preserved under addition. ∎
It is worth noting that Proposition 3.2 characterizes the class of upper-$\mathcal{C}^{2}$ functions as those satisfying a nonsmooth and local version of the descent lemma, given by (3). This property will be used to prove the convergence of the algorithmic schemes that we propose in Sections 4 and 5. In addition, Proposition 3.2 shows that the nonsmooth functions considered in [6, 8] are upper-$\mathcal{C}^{2}$ functions. Moreover, inequality (3) can be interpreted as a local counterpart of [8, Proposition 2.2]. Unlike that global result, Proposition 3.2 permits the constant $\kappa$ to vary with the point $\bar{x}\in U$. Notably, while the algorithm developed in [8] requires prior knowledge of $\kappa$, the subgradient methods that we shall introduce below do not rely on such information. Furthermore, Proposition 3.2 provides a more precise and localized version of [6, Lemma 3.6], where the descent property was established using a localization of the subgradient.
Let us present some examples that reveal that the class of upper-$\mathcal{C}^{2}$ functions naturally appears in several contexts in optimization. The first example illustrates the preservation of the notion under addition, multiplication by scalars, and composition.
@@
It is easy to prove that the class of upper-$\mathcal{C}^{2}$ is stable under addition and multiplication by positive scalars (see, e.g., [33, Exercise 10.35]). It is also closed under composition of upper-$\mathcal{C}^{2}$ and $\mathcal{C}^{1,+}$ functions. More precisely, given a function $\Phi$ that is $\mathcal{C}^{1,+}$ around $\bar{x}$, and an upper-$\mathcal{C}^{2}$ function $\varphi$ around $\Phi(\bar{x})$, then the function $\varphi\circ\Phi$ is upper-$\mathcal{C}^{2}$ around $\bar{x}$. Indeed, it follows from Proposition 3.2 that $\varphi(u)=\kappa\|u\|^{2}-h(u)$ locally around some neighborhood $V$ of $\Phi(\bar{x})$. Then, on some neighborhood of $\bar{x}$, one has $\varphi(\Phi(x))=\kappa\|\Phi(x)\|^{2}-h(\Phi(x))$, which is the difference of a differentiable function with locally Lipschitz gradient and a prox-regular function, since $h\circ\Phi$ is prox-regular due to [10, Proposition 2.3].
@@
First, let us notice that Proposition 3.2(iv) tells us that the difference of a $\mathcal{C}^{1,+}$ function and a continuous convex function (actually a prox-regular function) is indeed upper-$\mathcal{C}^{2}$. In the particular case where the differentiable function is also convex, this encompasses a class of difference of convex functions whose minimization has been addressed by several numerical methods in recent years (see, e.g., [17, 4, 5] and the references therein). Nonetheless, such works crucially rely on this decomposition to be available, which is not always easy to disclose (see, e.g., [22, 36]).
@@
This envelope plays a central role in mathematical programming and various branches of applied mathematics. Over the past few decades, its properties have been widely exploited in the development of numerical algorithms. In particular, it is well known that if $\psi$ is convex, then the Moreau envelope ${\mathtt{e}}_{\lambda}\psi$ is continuously differentiable and has a (globally) Lipschitz continuous gradient. More generally, when $\psi$ is not necessarily convex, it has been shown (see, e.g., [33, Example 10.32] or [6, Proposition 2.7]) that the Moreau envelope is an upper-$\mathcal{C}^{2}$ function provided that $\lambda\in(0,\lambda_{\psi})$, where
@@
is the so-called threshold of prox-boundedness of $\psi$. In more exact terms, for every $\lambda\in(0,\lambda_{\psi})$, the Moreau envelope, given by (4), admits the representation
@@
where $f\colon\mathbb{R}^{n}\to\mathbb{R}$ is $\mathcal{C}^{2,1}$ (twice continuously differentiable with a globally Lipschitz Hessian) with an $L_{f}$-Lipschitz continuous gradient, and $\psi\colon\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is a proper, extended-real-valued prox-bounded function with threshold $\lambda_{\psi}>0$. Patrinos and Bemporad introduced in [28] a new class of ad-hoc regularization techniques for handling problem (7), which is now commonly called the forward-backward envelope (see, e.g., [35] and the references therein). Precisely, the forward-backward envelope of the function $\varphi$ with parameter $\lambda>0$ is defined as
@@
A straightforward computation gives us the following representation of the augmented Lagrangian function
$\mathcal{L}(x,y,\lambda)=f(x)+{\mathtt{e}}_{\lambda}\psi(G(x)+\lambda^{-1}y)-\frac{1}{2\lambda}\|y\|^{2}.$
Therefore, due to the previous examples, the augmented Lagrangian function is indeed an upper-$\mathcal{C}^{2}$ function.
To finish this section, we note that upper-$\mathcal{C}^{2}$ functions also arise naturally in many applications, such as the minimum sum-of-squares clustering problem discussed in Sections 6 and 7.

### v2

@@ Abstract
In this paper we present a nonmonotone line search subgradient algorithm tailored to upper-$\mathcal{C}^{2}$ functions. This is a family of nonsmooth and nonconvex functions that satisfies a nonsmooth and local version of the descent lemma, making them suitable for line searches. We prove subsequential convergence of the proposed algorithm to a stationary point of the optimization problem. Our approach allows us to cover the setting of various subgradient algorithms, including Newton and quasi-Newton methods. In addition, we propose a specification of the general scheme, named Self-adaptive Nonmonotone Subgradient Method (SNSM), which automatically updates the parameters of the line search. Particular attention is paid to the minimum sum-of-squares clustering problem, for which we provide a concrete implementation of SNSM. We conclude with some numerical experiments where we exhibit the advantages of SNSM in comparison with some known algorithms.
@@ 1 Introduction
where the objective function $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ is assumed to be upper-$\mathcal{C}^{2}$. Roughly speaking, these functions admit a local representation as a minimum of twice-continuously differentiable functions (see Definition 3.1). Upper-$\mathcal{C}^{2}$ functions satisfy a nonsmooth local version of the descent lemma, a property that makes them suitable for extending the framework of line search methods (for details, see Section 3). This is important in practice, since upper-$\mathcal{C}^{2}$ functions naturally arise in different applications, as data mining (see Section 6).
Upper-$\mathcal{C}^{2}$ functions can equivalently be characterized as those that locally admit a difference of convex (DC) decomposition consisting of a smooth and a possibly nonsmooth convex functions (see Proposition 3.2). This might lead one to think that minimizing an upper-$\mathcal{C}^{2}$ function can be reduced to a standard DC programming task. However, this approach presents practical challenges. On the one hand, not all upper-$\mathcal{C}^{2}$ functions are globally DC (see Example 3.4). On the other hand, it requires prior knowledge of a suitable DC decomposition. For instance, given a DC function $f=g_{1}-g_{2}$ for some convex functions $g_{1}$ and $g_{2}$, it is often unclear which decomposition is optimal for numerical purposes, since $f=(g_{1}+h)-(g_{2}+h)$ yields another valid DC decomposition for any convex function $h$. This difficulty can be circumvented by treating the function $f$ as a unified entity, thereby eliminating the need for a specific DC decomposition.
Let us recall that descent line search methods are a class of iterative optimization algorithms designed to find (local) minima of differentiable functions. Their core technique consists in updating the current iterate by moving in a direction that reduces the value of the objective function. Hence, given a starting point $x_{0}\in\mathbb{R}^{n}$, a general iteration step is defined by
@@
While most DC algorithms only use first-order information, a coderivative-based semi-Newton method for difference programming has been recently proposed in [6]. When applied to a DC decomposition whose first component is twice differentiable, this method employs a Newton-type descent direction based on the Hessian of that component. However, since second-order information of the second function is not used, the method fails to capture the full curvature of the objective function. This limitation is evident in the minimum sum-of-squares clustering problem (see Section 6). For this problem, we demonstrate that effective second-order information can be successfully extracted and utilized by working directly with the upper-$\mathcal{C}^{2}$ function, which yields significantly faster algorithmic schemes (see Subsection 7.2). This further motivates our development of numerical schemes tailored to upper-$\mathcal{C}^{2}$ functions, avoiding the need to rely on DC decompositions.
@@
Although nonmonotone line searches bring opportunities to improve the performance of optimization algorithms, they also require the tuning of additional parameters, such as the line search stepsize (which is initially chosen at each iteration) and the memory parameter (which controls how many past function values are taken into account). A wrong adjustment of these parameters may slow down the resulting algorithm, and one may lose all possible benefits of nonmonotone line searches. With this potential weakness in mind, we propose a specification of the main scheme, that we name the Self-adaptive Nonmonotone Subgradient Method (SNSM), which incorporates an automatic procedure for selecting the stepsize and memory parameters of the line search. Other possibilities that would be interesting to investigate include the nonmonotone strategy proposed by Zhang and Hager [47], based on an average of function values.
The main contributions of this work are summarized as follows:
  • We characterize the variational properties that make upper-$\mathcal{C}^{2}$ functions suitable for optimization, allowing us to subsume existing approaches within a unified framework. Furthermore, we demonstrate their prevalence in optimization and provide examples to facilitate their identification.
  • We present a line search subgradient method tailored to upper-$\mathcal{C}^{2}$ functions (Algorithm 1), which permits the incorporation of second-order information. We prove its global subsequential convergence to stationary points of problem (P). Additionally, we propose a self-adaptive version of this general method (Algorithm 2) to mitigate the effects of suboptimal parameter configurations.
  • We conduct extensive numerical comparisons against specialized algorithms. First, we demonstrate how the nonmonotone line search generally leads to improved solutions for minimizing quadratic objectives with integer constraints. Second, we show that incorporating second-order information—which is typically lost in standard DC approaches—results in significantly superior algorithmic performance for the minimum sum-of-squares clustering problem. Overall, our experiments indicate that our method is competitive with modern large-scale solvers.
@@ 2 Preliminaries and notation
A function $f:\mathbb{R}^{n}\to\mathbb{R}$ is said to be of class $\mathcal{C}^{k}$ if it is differentiable and all its partial derivatives up to order $k$ are continuous. In particular, we say that a function $f$ is smooth if it is of class $\mathcal{C}^{1}$, i.e., it is differentiable with continuous gradient. If, in addition, its gradient is $L$-Lipschitz continuous we say that $f$ is $L$-smooth. The class of differentiable functions with a locally Lipschitz continuous gradient is denoted by $\mathcal{C}^{1,+}$. We represent the gradient and Hessian of $f$ as $\nabla f$ and $\nabla^{2}f$, respectively.
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a locally Lipschitz function around $\bar{x}$. We say that $f$ is prox-regular at a point point $\bar{x}$ if for every $\bar{v}\in\partial f(\bar{x})$ there exist $\varepsilon>0$ and $\rho>0$ such that
[…]
@@ 3 Upper-$\mathcal{C}^{2}$ functions and the descent lemma
Most known gradient-based algorithms are designed to ensure descent of the objective function per iteration. Many of the algorithms proposed in the literature that rely exclusively on first- and second-order information are developed for the setting of differentiable functions with at least locally Lipschitz continuous gradients. This class of functions is particularly important in both theoretical and practical optimization. Its significance lies in the fact that such functions satisfy the so-called descent lemma (see, e.g., [26, Lemma A.11]). More precisely, for a differentiable function $f:C\subseteq\mathbb{R}^{n}\to\mathbb{R}$ with an $L$-Lipschitz continuous gradient on a nonempty open convex set $C$, it holds that
$f(y)\leq f(x)+\langle\nabla f(x),y-x\rangle+\frac{L}{2}\|y-x\|^{2},\quad\text{for all }x,y\in C.$ (1)
Recall that $L$-smooth functions are those with global Lipschitz continuous gradients. Consequently, these functions verify the descent lemma (1) for $C=\mathbb{R}^{n}$. In contrast, for smooth functions whose gradients are only locally Lipschitz continuous on $\mathbb{R}^{n}$ (i.e., functions belonging to the class $\mathcal{C}^{1,+}$), the descent lemma may hold only locally around every point of the space. In such a case, for each $\bar{x}\in\mathbb{R}^{n}$, the descent lemma yields
$f(y)\leq f(x)+\langle\nabla f(x),y-x\rangle+\frac{L(\bar{x})}{2}\|y-x\|^{2},\quad\text{for all }x,y\in U(\bar{x}),$
where $U(\bar{x})$ is an open convex neighborhood of $\bar{x}$ and $L(\bar{x})\geq 0$ is a constant depending on both $\bar{x}$ and $U(\bar{x})$.
@@
The work [6] investigated objective functions that can be expressed as the difference of two functions, namely $\varphi:=g-h$, where $g$ is differentiable with a locally Lipschitz continuous gradient, and $h$ is a locally Lipschitz continuous prox-regular function (see, e.g., [38, Definition 13.27]). The authors developed a descent method for the objective function $\varphi$, where descent directions are generated using coderivative information. Among the results presented, they proved that this class of functions satisfies a local version of the descent lemma (see [6, Lemma 3.6]).
@@
Let us recall a related definition extracted from the book of Rockafellar and Wets (see [38, Definition 10.29]) that will serve us to encompass both frameworks [6] and [7]. It refers to a class of functions that is usually known as lower-$\mathcal{C}^{2}$ functions (correspondingly, a function $f$ is called upper-$\mathcal{C}^{2}$ if f-f is lower-$\mathcal{C}^{2}$). The properties of these functions have been well-studied in the context of variational analysis (see, e.g., [13, 34] and the references therein). However, to the best of our knowledge, this class has not received much attention in the numerical optimization literature. Beyond the already mentioned exceptions, we are only aware of [46], where a sequential quadratic programming trust-region method is proposed for the problem of minimizing a Lipschitz continuous and upper-$\mathcal{C}^{2}$ function in a bounded region including $L$-smooth equality constraints, and of [16], where the author studies a Frank–Wolfe method for nonsmooth upper-$\mathcal{C}^{1,\alpha}$ functions, relying on a generalized descent lemma for this class (see [16, Lemma 5]). It is also worth mentioning that, in [44, Chapter 14], the authors explore the structure of differences of weakly convex functions, consequently upper-$\mathcal{C}^{2}$, as upper-bounded models for nonsmooth optimization problems.
@@
It is easy to see that a function $\varphi$ is upper-$\mathcal{C}^{2}$ on an open set $U$ if and only if it is upper-$\mathcal{C}^{2}$ around any point $x\in U$. Moreover, by [38, Theorem 10.31], these functions are locally Lipschitz. The next result summarizes some useful characterizations of upper-$\mathcal{C}^{2}$ functions.
@@
where $g(x):=\kappa\|x\|^{2}$ and $h(x):=\max_{c\in C}\{\langle a(c),x\rangle+b(c)\}$. This shows the implication (ii) $\Rightarrow$ (iv). Finally, we conclude the proof by showing that (iv) $\Rightarrow$ (i). Let us suppose that there exists some neighborhood $V$ of $\bar{x}$ and a representation $\varphi=g-h$, where $g$ is differentiable with Lipschitz gradient on $V$ and $h$ is Lipschitz continuous and prox-regular on $V$. It follows from [38, Proposition 13.34] that $g$ is upper-$\mathcal{C}^{2}$ on $V$, and by [38, Proposition 13.33], the function h-h is upper-$\mathcal{C}^{2}$ on $V$. This concludes the proof, since the class of upper-$\mathcal{C}^{2}$ functions is preserved under addition. ∎
It is worth noting that Proposition 3.2 characterizes the class of upper-$\mathcal{C}^{2}$ functions as those satisfying a nonsmooth and local version of the descent lemma, given by (3). This property will be used to prove the convergence of the algorithmic schemes that we propose in Sections 4 and 5. In addition, Proposition 3.2 shows that the nonsmooth functions considered in [6, 7] are upper-$\mathcal{C}^{2}$ functions. Moreover, inequality (3) can be interpreted as a local counterpart of [7, Proposition 2.2]. Unlike that global result, Proposition 3.2 permits the constant $\kappa$ to vary with the point $\bar{x}\in U$. Notably, while the algorithm developed in [7] requires prior knowledge of $\kappa$, the subgradient methods that we shall introduce below do not rely on such information. Furthermore, Proposition 3.2 provides a more precise and localized version of [6, Lemma 3.6], where the descent property was established using a localization of the subgradient.
Let us present some examples that reveal that the class of upper-$\mathcal{C}^{2}$ functions naturally appears in several contexts in optimization. While there is no universal criterion to determine whether a function is upper-$\mathcal{C}^{2}$, the first two examples provide a useful toolkit for this purpose. Specifically, the first example illustrates the preservation of the property under addition, multiplication by scalars, and composition.
@@
It is easy to prove that the class of upper-$\mathcal{C}^{2}$ is stable under addition and multiplication by positive scalars (see, e.g., [38, Exercise 10.35]). It is also closed under composition of upper-$\mathcal{C}^{2}$ and $\mathcal{C}^{1,+}$ functions. More precisely, given a function $\Phi$ that is $\mathcal{C}^{1,+}$ around $\bar{x}$, and an upper-$\mathcal{C}^{2}$ function $\varphi$ around $\Phi(\bar{x})$, then the function $\varphi\circ\Phi$ is upper-$\mathcal{C}^{2}$ around $\bar{x}$. Indeed, it follows from Proposition 3.2 that $\varphi(u)=\kappa\|u\|^{2}-h(u)$ locally around some neighborhood $V$ of $\Phi(\bar{x})$. Then, on some neighborhood of $\bar{x}$, one has $\varphi(\Phi(x))=\kappa\|\Phi(x)\|^{2}-h(\Phi(x))$, which is the difference of a differentiable function with a locally Lipschitz gradient and a prox-regular function, since $h\circ\Phi$ is prox-regular due to [11, Proposition 2.3].
@@
First, let us notice that Proposition 3.2(iv) tells us that the difference of a $\mathcal{C}^{1,+}$ function and a continuous convex function (actually a prox-regular function) is indeed upper-$\mathcal{C}^{2}$. In the particular case where the differentiable function is also convex, this encompasses a class of difference of convex functions whose minimization has been addressed by several numerical methods in recent years (see, e.g., [20, 4, 5] and the references therein). Nonetheless, such works crucially rely on this decomposition to be available, which is not always easy to disclose (see, e.g., [25, 43]). Finally, let us emphasize that not every upper-$\mathcal{C}^{2}$ function is necessarily DC, since its domain need not be convex. For instance, the function $\varphi:\mathbb{R}\setminus\{0\}\to\mathbb{R}$ defined by
$\varphi(x)=x^{2}+\frac{1}{x^{2}}$
is $\mathcal{C}^{1,+}$ around every point of its domain, but it cannot be decomposed (globally) as the difference of two convex functions.
@@
This envelope plays a central role in mathematical programming and various branches of applied mathematics. Over the past few decades, its properties have been widely exploited in the development of numerical algorithms. In particular, it is well known that if $\psi$ is convex, then the Moreau envelope ${\mathtt{e}}_{\lambda}\psi$ is continuously differentiable and has a (globally) Lipschitz continuous gradient. More generally, when $\psi$ is not necessarily convex, it has been shown (see, e.g., [38, Example 10.32] or [6, Proposition 2.7]) that the Moreau envelope is an upper-$\mathcal{C}^{2}$ function provided that $\lambda\in(0,\lambda_{\psi})$, where
@@
is the so-called threshold of prox-boundedness of $\psi$. In more exact terms, for every $\lambda\in~\hskip-2.84526pt(0,\lambda_{\psi})$, the Moreau envelope, given by (4), admits the representation
@@
where $f\colon\mathbb{R}^{n}\to\mathbb{R}$ is a $\mathcal{C}^{2,1}$ function (twice continuously differentiable with a globally Lipschitz Hessian) with an $L_{f}$-Lipschitz continuous gradient, and $\psi\colon\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is a proper, extended-real-valued prox-bounded function with threshold $\lambda_{\psi}>0$. Patrinos and Bemporad introduced in [31] a new class of ad-hoc regularization techniques for handling problem (7), which is now commonly called the forward-backward envelope (see, e.g., [41] and the references therein). Precisely, the forward-backward envelope of the function $\varphi$ with parameter $\lambda>0$ is defined as
@@
In order to show that the augmented Lagrangian function is upper-$\mathcal{C}^{2}$, let us derive an alternative representation in terms of the Moreau envelope of $\psi$. By completing the squares on the terms $\frac{1}{2\lambda}\|u\|^{2}-\langle y,u\rangle$ and performing the change of variables $v:=G(x)+u$, we get
$\mathcal{L}(x,y,\lambda)$ $=\inf_{u\in\mathbb{R}^{m}}\left\{f(x)+\psi(G(x)+u)+\frac{1}{2\lambda}\|u-\lambda y\|^{2}-\frac{\lambda}{2}\|y\|^{2}\right\}$
$=f(x)+\inf_{u\in\mathbb{R}^{m}}\left\{\psi(G(x)+u)+\frac{1}{2\lambda}\|u-\lambda y\|^{2}\right\}-\frac{\lambda}{2}\|y\|^{2}$
$=f(x)+\inf_{v\in\mathbb{R}^{m}}\left\{\psi(v)+\frac{1}{2\lambda}\|v-(G(x)+\lambda y)\|^{2}\right\}-\frac{\lambda}{2}\|y\|^{2}$
$=f(x)+{\mathtt{e}}_{\lambda}\psi(G(x)+\lambda y)-\frac{\lambda}{2}\|y\|^{2}.$
Therefore, by the preceding examples, the augmented Lagrangian function is indeed an upper-$\mathcal{C}^{2}$ function.